The biochemistry that takes place inside cells depends on various elements, such as sodium, potassium, and calcium, that are dissolved in water as ions. These ions enter cells through narrow pores in the cell membrane known as ion channels. Each ion channel, which is formed from a specialized protein molecule, is selective for one type of ion. When an ion channel opens in a cell wall, monovalent (charge ee) ions flow through the channel at a rate of 1.0×107ions/s1.0×107ions/s. The potential difference across the ion channel is 70 mVmV. What is the power dissipation in the channel?
Answer #1

number of ions per sec = 10^7 ion / sec

current = I = number of ions per sec * charge of each ion

             I = 10^7 * 1.6 X 10^-19 = 1.6 X 10^-12 Amp


potential difference = V = 70 x 10^-3 V

power = p = V * I = 70 X 10^-3 * 1.6 X 10^-12 = 112 X 10^-15 Watt

             P = 112 x 10^-15 Watt